OpenStudy (fellowroot):

Find all solutions in C (complex) of the given equation. Z^3 =-27i

OpenStudy (anonymous):

z^3 + 27i = 0 (z-3i)(z^2 + 9+ 3zi) = 0 solve it u'll get 3 solutions..

OpenStudy (anonymous):

\[Z^3 = 27 * i^3\] Beacuse \[i^3 = -i\] \[Z = 3i\]

OpenStudy (anonymous):

@Fellowroot Did u Get it..)

OpenStudy (anonymous):

there r 3 solutions ..

OpenStudy (fellowroot):

Thanks.